A lon of 250$ is made for 9 months at a simple interest rate of 10% annum. What is the interest charge? What ammount is due after 9 month?


1
Expert's answer
2021-05-31T16:34:09-0400

"SI=P\times R\times T"

"=20\times\frac{10}{100}\times \frac{9}{12}"

"=\$18.75"

"Amount =Interest+Principal\\=18.75+250\\=\$268.75"